# DAWN-SI: Data-aware and Noise-Informed Stochastic Interpolation for Solving Inverse Problems

### v1

@@ Abstract
Inverse problems, which involve estimating parameters from incomplete or noisy observations, arise in various fields such as medical imaging, geophysics, and signal processing. These problems are often ill-posed, requiring regularization techniques to stabilize the solution. In this work, we employ Stochastic Interpolation (SI), a generative framework that integrates both deterministic and stochastic processes to map a simple reference distribution, such as a Gaussian, to the target distribution. Our method DAWN-SI: Data-AWare and Noise-informed Stochastic Interpolation incorporates data and noise embedding, allowing the model to access representations about the measured data explicitly and also account for noise in the observations, making it particularly robust in scenarios where data is noisy or incomplete. By learning a time-dependent velocity field, SI not only provides accurate solutions but also enables uncertainty quantification by generating multiple plausible outcomes. Unlike pre-trained diffusion models, which may struggle in highly ill-posed settings, our approach is trained specifically for each inverse problem and adapts to varying noise levels. We validate the effectiveness and robustness of our method through extensive numerical experiments on tasks such as image deblurring and tomography.
@@ 1 Introduction
Figure 1: Example images from MNIST (left) and STL10 (right) datasets showing the recovery of deblurred images from the blurred (noisy) data (top panel) at different levels of noise (0%, 1%, 5%, 10%, 20%). Our methods DAW-SI  (middle panel) incorporates the embedding for the blurred data within the network, while DAWN-SI (bottom panel) incorporates embedding for both data and noise levels. DAWN-SI is superior to DAW-SI at deblurred image recovery, especially at higher levels of noise.
@@
Inverse problems are a class of problems in which the goal is to determine parameters (or parameter function) of a system from observed data. These problems arise in various fields, including medical imaging, geophysics, remote sensing, and signal processing. Inverse problems are often ill-posed, meaning that a unique solution does not exist, or the solution may be highly sensitive to small perturbations in the data. Solving inverse problems typically requires regularization techniques, which introduce additional constraints or prior information to stabilize the solution and mitigate the effects of ill-posedness. Such regularization can be obtained using various mathematical and computational methods, including optimization techniques (variational methods), statistical inference (Bayesian or frequentist) and machine learning [Scherzer et al., 2009, Adler and Öktem, 2017, Calvetti and Somersalo, 2024, López-Tapia et al., 2021]. In this paper, we approach inverse problems using the latter and investigate a new set of methods that are machine learning based for the solution of inverse problems. In particular, we show how stochastic interpolation (SI) which was recently proposed in the context of generative models can be effectively used to estimate the solution of inverse problems and to further investigate its non-uniqueness.
@@
Stochastic Interpolation is a relatively new generative process that provides a unifying framework, elegantly integrating both deterministic flows and stochastic diffusion models [Albergo et al., 2023]. The core concept of SI is to learn a stochastic process that effectively transports a simple reference distribution, such as Gaussian, to the desired target data distribution. This transportation process can manifest as either deterministic or stochastic. In the former case, it is described by an ordinary differential equation (ODE), while in the latter, it is governed by a stochastic differential equation (SDE).
@@
The stochastic interpolation framework defines a continuous-time reversibility between the reference and target distributions, parameterized by time $t\in[0,1]$. At the initial time $t=0$, the distribution aligns with the reference distribution. As time progresses to $t=1$, the distribution evolves to match the target data distribution. This evolution is achieved by learning the time-dependent velocity field (for ODEs) or drift and diffusion coefficients (for SDEs) that characterize this interpolation process. By understanding and modeling this time-dependent transformation, one can generate samples from the target distribution through numerical integration of the learned ODE or SDE.
@@
SI is a highly flexible methodology for designing new types of generative models. In this work, we use this flexibility and show how to adopt SI to solve highly ill-posed inverse problems. We find SI particularly useful since it allows ease of sampling from the target distribution. This implies that we are able to generate a range of solutions to the inverse problems and thus investigate the uncertainty that is associated with the estimated solution. Such a process is highly important in physical applications when decisions are made based on the solution.
Related work:
@@
The methods proposed here are closely related to three different approaches for the solution of inverse problems. First, there is an obvious link to the incorporation of diffusion models as regularizers in inverse problems [Yang et al., 2022, Chung et al., 2022a, c, b, Song et al., 2022]. The key idea is to leverage a pre-trained diffusion model that captures the data distribution as a prior, and then condition this model on the given measurements to infer the underlying clean signal or image. For inverse problems, diffusion methods condition the diffusion model on the given measurements (e.g. noisy, incomplete, or compressed data) by incorporating them into the denoising process. Nonetheless, it has been shown in Eliasof et al. [2024] that pre-trained diffusion models that are used for ill-posed inverse problems as regularizers tend to under-perform as compared to the models that are trained specifically on a particular inverse problem. In particular, such models tend to break when the noise level is not very low.
@@
Main Contribution:
The core contributions of this paper lie in the design and application of the DAWN-SI framework, specifically crafted for solving highly ill-posed inverse problems like image deblurring or tomography. (i) Our method is designed to be problem-specific, which adjusts itself to the unique structure of the inverse problem it is tasked with, learning the posterior distribution directly, ensuring that the learned velocity fields and mappings are directly applicable to the target task. (ii) We train a stochastic interpolant by embedding measured data and noise information directly into the interpolation process. This incorporation allows our model to adapt to a wide range of noise conditions, something that pre-trained models struggle with. By making the training explicitly aware of the noise level and data characteristics, our model can better navigate noisy or incomplete measured data, producing superior reconstructions. (iii) By learning the posterior distribution directly and leveraging the stochastic nature of the interpolation process, DAWN-SI generates multiple plausible solutions for a given inverse problem, allowing us to explore the solution space more thoroughly and in particular, to estimate the posterior mean and its standard deviation, estimating the uncertainty in the recovered solution.
2 Stochastic Interpolation and Inverse Problems
In this section we review stochastic interpolation as well as derive the main ideas behind using it for the solution of inverse problems.
2.1 Stochastic Interpolation: A Partial Review
Stochastic interpolation (SI) is a framework that transforms points between two distributions. Given two densities $\pi_{0}({\bf x})$ and $\pi_{1}({\bf x})$, the goal is to find a mapping that takes a point ${\bf x}_{0}\sim\pi_{0}({\bf x})$ and transports it to a point ${\bf x}_{1}\sim\pi_{1}({\bf x})$. For simplicity and for the purpose of this work, we choose $\pi_{0}$ to be a Gaussian distribution with 00 mean and ${\bf I}$ covariance.
@@ 2.1 Stochastic Interpolation: A Partial Review
Under the SI framework, the velocity field for all $({\bf x}_{t},t)$ is learned by averaging over all possible paths. To this end, we parameterize the velocity by a function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ and solve the stochastic optimization problem for ${\boldsymbol{\theta}}$
@@
Equation 5 is useful when a recovered ${\bf x}_{1}$ is desirable. In the context of using SI for inverse problems, obtaining an approximation to ${\bf x}_{1}$ can be desirable, as we see next.
2.2 Applying Stochastic Interpolation for the solution of inverse problems
@@
If ${\boldsymbol{\lambda}}_{i}$ is large, the projection of ${\bf x}_{1}$ onto the eigenvector ${\bf v}_{i}$ is very informative and very minimal regularization is required. However, if ${\boldsymbol{\lambda}}_{i}\approx 0$, the contribution of ${\bf v}_{i}$ to the solution is difficult, if not impossible, to obtain and this is where the regularization is highly needed. When using pre-trained models, the prior $\pi({\bf x}_{1})$ is estimated numerically and it is unaware of the inverse problem at hand. Errors in the estimated $\pi({\bf x}_{1})$ in the parts that correspond to the large singular vectors may not be destructive. However, if $\pi({\bf x}_{1})$ has errors that correspond to the very small singular values, that is, to the effective null space of the data, this may lead to the artifacts that have been observed early in Kaipio and Somersalo [2004], Tenorio et al. [2011]. This suggests that although it is appealing to use a generic pre-trained priors in the process, a better approach is to not use the Bayesian factorization to prior and likelihood but rather train a stochastic interpolant that maps the distribution $\pi({\bf x}_{0})$ to the posterior $\pi({\bf x}_{1}|{\bf b})$ directly. Indeed, as we show next, the flexibility of the SI framework allows us to learn a velocity function that achieves just that.
@@ 2.3 A data-aware and noise-informed velocity estimator
In the canonical form of stochastic interpolation, the velocity is estimated from the interpolated vector ${\bf x}_{t}$. In the context of inverse problems, we also have a vector ${\bf b}$ (measured data) that contains additional information on ${\bf x}_{1}$ and therefore can be used to estimate the velocity towards the posterior. We now show that by doing a small change to the training process of SI, it is possible to solve inverse problems using the same concept.
@@
An important question is the design of a network that integrates the information about ${\bf b}$ into the velocity estimation process. One important choice is the function $f$ that operates on ${\bf b}$. For many, if not most, inverse problems, the data ${\bf b}$ belongs to a different space than ${\bf x}$. Therefore, it is difficult to use this vector directly. The goal of the function $f$ is to transform the data, ${\bf b}$ from the data space to the space of ${\bf x}$. One obvious approach to achieve this is to choose
@@
Assume that we have a prior knowledge that the distribution of ${\bf x}$ is composed of two Gaussians. Using SI to generate data from these Gaussians is demonstrated in Figure 2.
@@
Figure 2: Generation of the distribution of two Gaussians by SI without and with data.
Figure 3: Schematic of the training process for the SI model for solving inverse problem, where the forward model is given by ${\bf A}$. This figure specifically represents the deblurring inverse problem. The network with parameters ${\boldsymbol{\theta}}$ can either have noise embedding (DAWN-SI) or not (DAW-SI). The two loss terms $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the error in prediction of velocity and misfit, respectively. In this figure, the transformation $f$ for generating the data embedding was chosen as $f={\bf A}^{\top}$.
2.4 Training the SI model for inverse problems
When training an SI model, we solve the optimization problem given by Equation 3. In this process, one draws samples from ${\bf x}_{0}$ and ${\bf x}_{1}$, then randomly chooses $t\in[0,1]$ to generate the vector ${\bf x}_{t}$. In addition, we generate the data ${\bf b}$ by multiplying the matrix ${\bf A}$ with ${\bf x}_{1}$ and adding noise $\epsilon$ with a random standard deviation $\sigma$.
Next, one feeds the network ${\bf x}_{t}$, $t$, ${\bf b}$ and $\sigma$ to compute ${\bf s}_{\theta}({\bf x}_{t},{\bf b},t,\sigma)$ and then compare it to the velocity ${\bf v}={\bf x}_{1}-{\bf x}_{0}$.
While it is possible to train the network by comparing the computed velocity to its theoretical value, we have found that adding additional terms to the loss helps in getting a better estimate for the velocity. In particular, when training the model for inverse problems, we add an additional term to the loss that relates to the particular inverse problem in mind. To this end, note that after estimating ${\bf v}$, we can estimate ${\bf x}_{1}$ using Equation 5. Thus, we can estimate the data fit ${\bf b}_{{\boldsymbol{\theta}}}$ for the estimated velocity, given by,
@@ 2.4 Training the SI model for inverse problems
where $\alpha$ is a hyperparameter, which we set to 1. A schematic for training of SI model is given in Figure 3.
@@ 2.5 Architectures for data-aware and noise-informed velocity estimators
The embedding of time and noise is straight-forward. Note that both $t$ and $\sigma$ are scalars. For their embedding, we use the method presented in Croitoru et al. [2023], which involves creating learnable embedding and adding them to the feature maps at each level of the network.
A key component in making SI perform well for inverse problems is the careful design of a network that incorporates the data within the training process. The resulting network can be thought of as a likelihood-free estimator [Thomas et al., 2022], that estimates the velocity vector ${\bf v}$, given the input vector ${\bf x}_{t}$, the time $t$ and the noise level $\sigma$. As previously explained, we do not integrate ${\bf b}$ directly, but rather use $f({\bf b})={\bf A}^{\top}{\bf b}$ and embed this vector in the network.
@@
be an encoder network that is parameterized by ${\boldsymbol{\eta}}$. The encoder can be pre-trained or trained as part of the network, providing flexibility in how it is integrated into the UNet. The encoded data provides additional context that is critical for accurate predictions. Formally, the UNet estimates the velocity as follows:
@@ 2.6 Inference and Uncertainty Estimation
Figure 4: Computing uncertainty in the solutions obtained by SI for some example images from MNIST (left), STL10 (middle) and CIFAR10 (right) for the deblurring task. The reconstruction (posterior mean) is computed by averaging over runs from 32 randomly initialized ${\bf x}_{0}$. The uncertainty in the posterior mean is computed by evaluating its standard deviation for the predictions from 32 runs.
@@
An example for this process is shown in Figure 4. Here, the solutions over multiple runs were averaged to generate an effective reconstruction of the original image (the posterior mean). Moreover, to quantify uncertainty in the reconstruction process, we computed standard deviation over the solutions at each pixel from multiple runs. In the uncertainty maps for MNIST, the uncertainty is concentrated along the edges of the digits. This occurs because the SI model introduces slight variations in how it reconstructs the boundary between the digit and the background. Since this boundary is sharp, any slight differences in how this edge is defined in different reconstructions lead to higher uncertainty along the edges. On the other hand, for STL10 and CIFAR10 images, the boundary between objects and background is often less distinct. The background might contain detailed textures or noise that blends into the object, making it harder for the model to distinguish clear boundaries. Hence, the uncertainty maps for these datasets do not exhibit the same clear edge-focused uncertainty as in MNIST. The lack of a clear boundary means that the reconstruction’s variability spreads more evenly across the entire image.

### v2

# DAWN-FM: Data-Aware and Noise-Informed Flow Matching for Solving Inverse Problems

@@ Abstract
Inverse problems, which involve estimating parameters from incomplete or noisy observations, arise in various fields such as medical imaging, geophysics, and signal processing. These problems are often ill-posed, requiring regularization techniques to stabilize the solution. In this work, we employ Flow Matching (FM), a generative framework that integrates a deterministic processes to map a simple reference distribution, such as a Gaussian, to the target distribution. Our method DAWN-FM: Data-AWare and Noise-informed Flow Matching incorporates data and noise embedding, allowing the model to access representations about the measured data explicitly and also account for noise in the observations, making it particularly robust in scenarios where data is noisy or incomplete. By learning a time-dependent velocity field, FM not only provides accurate solutions but also enables uncertainty quantification by generating multiple plausible outcomes. Unlike pre-trained diffusion models, which may struggle in highly ill-posed settings, our approach is trained specifically for each inverse problem and adapts to varying noise levels. We validate the effectiveness and robustness of our method through extensive numerical experiments on tasks such as image deblurring and tomography.
@@ 1 Introduction
Figure 1: Example images from MNIST (left) and STL10 (right) datasets showing the recovery of deblurred images from the blurred (noisy) data (top panel) at different levels of noise (0%, 1%, 5%, 10%, 20%). Our methods DAW-FM  (middle panel) incorporates the embedding for the blurred data within the network, while DAWN-FM (bottom panel) incorporates embedding for both data and noise levels. DAWN-FM is superior to DAW-FM at deblurred image recovery, especially at higher levels of noise.
@@
Inverse problems are a class of problems in which the goal is to determine parameters (or parameter function) of a system from observed data. These problems arise in various fields, including medical imaging, geophysics, remote sensing, and signal processing. Inverse problems are often ill-posed, meaning that a unique solution does not exist, or the solution may be highly sensitive to small perturbations in the data. Solving inverse problems typically requires regularization techniques, which introduce additional constraints or prior information to stabilize the solution and mitigate the effects of ill-posedness. Such regularization can be obtained using various mathematical and computational methods, including optimization techniques (variational methods), statistical inference (Bayesian or frequentist) and machine learning [Scherzer et al., 2009, Adler and Öktem, 2017, Calvetti and Somersalo, 2024, López-Tapia et al., 2021]. In this paper, we approach inverse problems using the latter and investigate a new set of methods that are machine learning based for the solution of inverse problems. In particular, we show how flow matching (FM) which was recently proposed in the context of generative models can be effectively used to estimate the solution of inverse problems and to further investigate its non-uniqueness.
@@
Stochastic Interpolation (SI) is a relatively new generative process that provides a unifying framework, elegantly integrating both deterministic flows and stochastic diffusion models [Albergo and Vanden-Eijnden, 2022, Albergo et al., 2023]. The core concept of SI is to learn a stochastic process that effectively transports a simple reference distribution, such as Gaussian, to the desired target data distribution. This transportation process can manifest as either deterministic or stochastic. In the former case, it is described by an ordinary differential equation (ODE), while in the latter, it is governed by a stochastic differential equation (SDE). For this work, we rely only on the formal deterministic ODE formalism of SI, which is known as Flow Matching (FM).
@@
The FM framework defines a continuous-time reversibility between the reference and target distributions, parameterized by time $t\in[0,1]$. At the initial time $t=0$, the distribution aligns with the reference distribution. As time progresses to $t=1$, the distribution evolves to match the target data distribution. This evolution is achieved by learning the time-dependent velocity (or drift) field that characterize this interpolation process. By understanding and modeling this time-dependent transformation, one can generate samples from the target distribution through numerical integration of the learned ODE.
@@
FM is a highly flexible methodology for designing new types of generative models. In this work, we use this flexibility and show how to adopt FM to solve highly ill-posed inverse problems. We find FM particularly useful since it allows ease of sampling from the target distribution. This implies that we are able to generate a range of solutions to the inverse problems and thus investigate the uncertainty that is associated with the estimated solution. Such a process is highly important in physical applications when decisions are made based on the solution.
@@
Related work: The methods proposed here are closely related to three different approaches for the solution of inverse problems. First, there is an obvious link to the incorporation of diffusion models as regularizers in inverse problems [Yang et al., 2022, Chung et al., 2022a, c, b, Song et al., 2022]. The key idea is to leverage a pre-trained diffusion model that captures the data distribution as a prior, and then condition this model on the given measurements to infer the underlying clean signal or image. For inverse problems, diffusion methods condition the diffusion model on the given measurements (e.g. noisy, incomplete, or compressed data) by incorporating them into the denoising process. Nonetheless, it has been shown in Eliasof et al. [2024] that pre-trained diffusion models that are used for ill-posed inverse problems as regularizers tend to under-perform as compared to the models that are trained specifically on a particular inverse problem. In particular, such models tend to break when the noise level is not very low.
@@
Main contribution: The core contributions of this paper lie in the design and application of the DAWN-FM framework, specifically crafted for solving highly ill-posed inverse problems like image deblurring and tomography. (i) Our method is designed to be problem-specific, which adjusts itself to the unique structure of the inverse problem it is tasked with, learning the posterior distribution directly, ensuring that the learned velocity fields and mappings are directly applicable to the target task. (ii) We train an interpolant by embedding measured data and noise information directly into the interpolation process. This incorporation allows our model to adapt to a wide range of noise conditions, something that pre-trained models struggle with. By making the training explicitly aware of the noise level and data characteristics, our model can better navigate noisy or incomplete measured data, producing superior reconstructions. (iii) By learning the posterior distribution directly and leveraging the stochastic nature of the interpolation process, DAWN-FM generates multiple plausible solutions for a given inverse problem, allowing us to explore the solution space more thoroughly and in particular, to estimate the posterior mean and its standard deviation, estimating the uncertainty in the recovered solution.
2 Flow Matching and Inverse Problems
In this section we review flow matching as well as derive the main ideas behind using it for the solution of inverse problems.
2.1 Flow Matching (FM): A partial review
Flow Matching (FM) is a framework that transforms points between two distributions. Given two densities $\pi_{0}({\bf x})$ and $\pi_{1}({\bf x})$, the goal is to find a mapping that takes a point ${\bf x}_{0}\sim\pi_{0}({\bf x})$ and transports it to a point ${\bf x}_{1}\sim\pi_{1}({\bf x})$. For simplicity and for the purpose of this work, we choose $\pi_{0}$ to be a Gaussian distribution with 00 mean and ${\bf I}$ covariance.
@@ 2.1 Flow Matching (FM): A partial review
Under the FM framework, the velocity field for all $({\bf x}_{t},t)$ is learned by averaging over all possible paths. To this end, we parameterize the velocity by a function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ and solve the optimization problem for ${\boldsymbol{\theta}}$,
@@
Equation 5 is useful when a recovered ${\bf x}_{1}$ is desirable. In the context of using FM for inverse problems, obtaining an approximation to ${\bf x}_{1}$ can be desirable, as we see next.
2.2 Applying FM for solving inverse problems
@@
If ${\boldsymbol{\lambda}}_{i}$ is large, the projection of ${\bf x}_{1}$ onto the eigenvector ${\bf v}_{i}$ is very informative and very minimal regularization is required. However, if ${\boldsymbol{\lambda}}_{i}\approx 0$, the contribution of ${\bf v}_{i}$ to the solution is difficult, if not impossible, to obtain and this is where the regularization is highly needed. When using pre-trained models, the prior $\pi({\bf x}_{1})$ is estimated numerically and it is unaware of the inverse problem at hand. Errors in the estimated $\pi({\bf x}_{1})$ in the parts that correspond to the large singular vectors may not be destructive. However, if $\pi({\bf x}_{1})$ has errors that correspond to the very small singular values, that is, to the effective null space of the data, this may lead to the artifacts that have been observed early in Kaipio and Somersalo [2004], Tenorio et al. [2011]. This suggests that, although it is appealing to use a generic pre-trained priors in the process, a better approach is to not use the Bayesian factorization to prior and likelihood but rather train a flow matching interpolant that maps the distribution $\pi({\bf x}_{0})$ to the posterior $\pi({\bf x}_{1}|{\bf b})$ directly. Indeed, as we show next, the flexibility of the FM framework allows us to learn a velocity function that achieves just that.
@@ 2.3 A data-aware and noise-informed velocity estimator
In the canonical form of FM, the velocity is estimated from the interpolated vector ${\bf x}_{t}$. In the context of inverse problems, we also have a vector ${\bf b}$ (measured data) that contains additional information on ${\bf x}_{1}$ and therefore can be used to estimate the velocity towards the posterior. We now show that by doing a small change to the training process of FM, it is possible to solve inverse problems using the same concept.
@@
An important question is the design of a network that integrates the information about ${\bf b}$ into the velocity estimation process. One important choice is the function $f$ that operates on ${\bf b}$. For many, if not most, inverse problems, the data ${\bf b}$ belongs to a different space than ${\bf x}$. Therefore, it is difficult to use this vector directly. The goal of the function $f$ is to transform ${\bf b}$ from the data space to the space of ${\bf x}$. One obvious approach to achieve this is to choose
@@
Assume that we have a prior knowledge that the distribution of ${\bf x}$ is composed of two Gaussians. Using FM to generate data from these Gaussians is demonstrated in Figure 2.
@@
Figure 2: Generation of the distribution of two Gaussians by FM without and with data.
Figure 3: Schematic of the training process for the FM model for solving inverse problem, where the forward model is given by ${\bf A}$. This figure specifically represents the deblurring inverse problem. The network with parameters ${\boldsymbol{\theta}}$ can either have noise embedding (DAWN-FM) or not (DAW-FM). The two loss terms $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the error in prediction of velocity and misfit, respectively. In this figure, the transformation $f$ for generating the data embedding was chosen as $f={\bf A}^{\top}$.
2.4 Training the FM model for inverse problems
When training the FM model, we solve the optimization problem given by Equation 3. In this process, one draws samples from ${\bf x}_{0}$ and ${\bf x}_{1}$, then randomly chooses $t\in[0,1]$ to generate the vector ${\bf x}_{t}$. In addition, we generate the data ${\bf b}$ by multiplying the matrix ${\bf A}$ with ${\bf x}_{1}$ and adding noise $\epsilon$ with a random standard deviation $\sigma$.
Next, one feeds the network ${\bf x}_{t}$, $t$, ${\bf b}$ and $\sigma$ to compute ${\bf s}_{\theta}({\bf x}_{t},{\bf b},t,\sigma)$ and then compare it to the true velocity ${\bf v}={\bf x}_{1}-{\bf x}_{0}$.
While it is possible to train the network by comparing the computed velocity to its theoretical value, we have found that adding additional terms to the loss helps in getting a better estimate for the true velocity. In particular, when training the model for inverse problems, we add an additional term to the loss that relates to the particular inverse problem in mind. To this end, note that after estimating ${\bf v}$, we can estimate ${\bf x}_{1}$ using Equation 5. Thus, we can estimate the data fit ${\bf b}_{{\boldsymbol{\theta}}}$ for the estimated velocity, given by,
@@ 2.4 Training the FM model for inverse problems
where $\alpha$ is a hyperparameter, which we set to 1. A schematic for training of FM model is given in Figure 3.
@@ 2.5 Architectures for data-aware and noise-informed velocity estimators
The embedding of time and noise is straightforward. Note that both $t$ and $\sigma$ are scalars. For their embedding, we use the method presented in Croitoru et al. [2023], which involves creating learnable embedding and adding them to the feature maps at each level of the network.
A key component in making FM perform well for inverse problems is the careful design of a network that incorporates the data within the training process. The resulting network can be thought of as a likelihood-free estimator [Thomas et al., 2022], that estimates the velocity vector ${\bf v}$, given the input vector ${\bf x}_{t}$, the time $t$ and the noise level $\sigma$. As previously explained, we do not integrate ${\bf b}$ directly, but rather use $f({\bf b})={\bf A}^{\top}{\bf b}$ and embed this vector in the network.
@@
be an encoder network that is parameterized by ${\boldsymbol{\eta}}$. The encoder can be pre-trained or trained as part of the network, providing flexibility in how it is integrated into the UNet. The encoded data provides additional context that is critical for accurate predictions. Formally, the UNet estimates the velocity as,
@@ 2.6 Inference and uncertainty estimation
Figure 4: Computing uncertainty in the solutions obtained by FM for some example images from MNIST (left), STL10 (middle) and CIFAR10 (right) for the deblurring task. The reconstruction (posterior mean) is computed by averaging over runs from 32 randomly initialized ${\bf x}_{0}$. The uncertainty in the posterior mean is computed by evaluating its standard deviation for the predictions from 32 runs.
@@
An example for this process is shown in Figure 4. Here, the solutions over multiple runs were averaged to generate an effective reconstruction of the original image (the posterior mean). Moreover, to quantify uncertainty in the reconstruction process, we computed standard deviation over the solutions at each pixel from multiple runs. In the uncertainty maps for MNIST, the uncertainty is concentrated along the edges of the digits. This occurs because the FM model introduces slight variations in how it reconstructs the boundary between the digit and the background. Since this boundary is sharp, any slight differences in how this edge is defined in different reconstructions lead to higher uncertainty along the edges. On the other hand, for STL10 and CIFAR10 images, the boundary between objects and background is often less distinct. The background might contain detailed textures or noise that blends into the object, making it harder for the model to distinguish clear boundaries. Hence, the uncertainty maps for these datasets do not exhibit the same clear edge-focused uncertainty as in MNIST. The lack of a clear boundary means that the reconstruction’s variability spreads more evenly across the entire image.